Dunstreet's Department Store would like to develop an inventory ordering policy with a 99 percent probability of not stocking out. To illustrate your recommended procedure, use as an example the ordering policy for white percale sheets.
Demand for white percale sheets is 4,300 per year. The store is open 365 days per year. Every four weeks (28 days) inventory is counted and a new order is placed. It takes 12 days for the sheets to be delivered. Standard deviation of demand for the sheets is eight per day. There are currently 160 sheets on-hand.

a. How many sheets should you order? (Use Excel's NORM.S.INV() function or z-table to find the z value. Do not round intermediate calculations. Round z value to 2 decimal places and final answer to the nearest whole number)

Multi-period Model- Fixed-Order Quantity Model (Q-model)
Answer #1

Information available is as under:

D = Days operational

365

days

A = Annual Demand

4300

sheets per year

d = Average Demand rate per time period

11.78

per day

σ = Standard Deviation

8

bags

LT = Performance cycle, another term for total lead time

12

days

z for 99% target service level = NORM.S.INV(0.99) =

2.33

R = Reorder period or Review Period

28

days

T1 = time increment used for calculating standard deviation of demand

1

days

Stock on hand (current at the time of review)

160

Sheets

Answer a:

Restocking level = (d*(R+LT))+(z*(σ*(R+LT)) =

1217.833

sheets

Number of sheets to order =

Restocking level – Stock on hand = 1217.833 -160 = 1056.833 ≈

1057

sheets

Therefore, 1057 sheets needs to be ordered at this review period.